# TOMG-Bench: Evaluating LLMs on Text-based Open Molecule Generation

### v1

@@ 5 Conclusion
In this study, we introduce TOMG-Bench, the first benchmark designed to assess the capabilities of Large Language Models (LLMs) in the realm of open-domain molecule generation. Benchmarking 25 LLMs, TOMG-Bench highlights the limitations of existing targeted molecule generation tasks and demonstrates the potential of general LLMs in this domain. Additionally, through instruction tuning on our proposed OpenMolIns, LLMs exhibit significant potential on the TOMG-Bench, matching the performance of GPT-3.5-turbo. Our contributions not only lie in the development of a novel benchmark for molecular discovery but also provide a diverse indication of the capabilities of LLMs.
6 Limitations
Although TOMG-Bench is carefully designed and well-validated through our experiments, we still observe several limitations:
Prompt Diversity. Prompt diversity helps relieve the over-fitting of the instructions. While we adopt several different prompt templates and randomly choose from them, we still find the number of prompt templates is not enough to satisfy the prompt diversity.
Data Distribution. In our data construction process, we allocate distributions to atoms, bonds, and functional groups with the aim of making our benchmark more reflective of real-world distributions. Nevertheless, the distribution we use is largely empirical and may not be sufficiently accurate to reconstruct real-world scenarios accurately. This could potentially mask the true performance capabilities of LLMs in these specific tasks.
7 Acknowledgments
We thank all the reviewers for their insightful comments. We also thank GitHub Copilot for coding assistance and ChatGLM-4 for polishing the writing.
Appendix A Data Construction
In this section, we introduce the construction details of TOMG-Bench and OpenMolIns dataset, as well as the prompt templates.
A.1 MolEdit
@@
For the molecule editing (MolEdit) task, we consider the common operations on modifying functional groups in a given molecule (i.e., add, drop, and substitute), which are simple tasks for human experts but challenging to LLMs. In this case, we further develop three corresponding subtasks: AddComponent, DelComponent, and SubComponent. Prompt templates for MolEdit are shown in Table 2. However, there are different kinds of functional groups, and some functional groups can play an important role in the molecule structure, such as connecting two separate parts of the molecule, which makes them unsuitable for these operations above as these operations will entirely change the structure of the molecule. In this case, we aim to make a slight change in the molecule structure and limit most of the functional groups we choose within the end groups.
@@
Table 2: Prompt Templates for MolEdit
Table 3 presents the functional groups that are taken into account for AddComponent and DelComponent, along with their respective selection weights. To reflect the distribution of these functional groups in real-world scenarios, we have implemented a weighted random selection process for AddComponent, which ensures that less common functional groups are assigned a lower probability to be chosen, thereby refining the selection mechanism to better mirror practical occurrences.
@@
For SubComponent, our focus is exclusively on end groups, which include hydroxyl, aldehyde, carboxyl, nitro, halo, nitrile, and thiol, which ensures that the editing operations are confined to substituting the existing functional group with another from this list, thereby maintaining the integrity of the molecule’s overall structure without altering it fundamentally.
A.2 MolOpt
Molecule optimization (MolOpt), designed to optimize molecular properties through the refinement of molecule structures, is not a brand-new task. Previously, GNN-based methods have been widely adopted in this task, while these methods can only help with one specific subtask at a time. In contrast, TOMG-Bench requires one single LLM to optimize molecules with different metrics and directions. In this work, we specifically focus on enhancing specific characteristics that are crucial for drug discovery and chemical synthesis, including LogP, MR, and QED. The prompt templates for MolOpt are illustrated in Table 4.
@@ A.2 MolOpt
Table 4: Prompt Templates for MolOpt
@@
A.3 MolCustom
To enable customized design of molecules, we think of three fundamental features for describing the molecule, including atoms, bonds, and functional groups. Given the specified category and number of atoms, bonds, and functional groups, LLMs should generate the molecule as we request. The prompt templates for MolCustom are shown in Table 5. Below, we present the construction details of the three subtasks for MolCustom:
@@ A.3 MolCustom
Table 5: Prompt Templates for MolCustom
@@
Table 6: Atoms that are considered in AtomNum, as well as their weights to be selected.
@@
Table 7: Chemical bonds that are considered in BondNum, as well as their weights to be selected.
@@
Table 8: Functional Groups that are considered in FunctionalGroup, as well as their weights to be selected.
@@
Notably, in MolCustom, if not specified, LLMs can generate any number of these atoms, bonds, and functional groups. However, for these specified items, LLMs should strictly follow the requirements.
Appendix B Hyper Parameters
Item Value
Generation
temperature 0.75
top_p 0.85
num_beams 1
max_new_tokens 512
Instruction Tuning
batchsize 32
lr 3e-4
cutoff_len 1024
Lora Settings
r 64
$\alpha$ 128
dropout 0.1
Table 9: Hyper-parameters
Appendix C Detailed Results
@@
In this section, we first show the leaderboard of TOMG-Bench in Table 10, where Claude-3.5 achieves first place with a weighted average accuracy of 35.92%. Notably, via instruction tuning on our OpenMolIns dataset, Llama-3.1-8B achieves 6th place, which outperforms all the existing open-source LLMs and is just behind Cladue-3.
Model #Parameters (B) $\bar{Acc}$ (%) $\bar{wAcc}(\%)$
Claude-3.5 Anthropic (2024b) - 51.10 35.92
Gemini-1.5-pro Deepmind (2024) - 52.25 34.80
GPT-4-turbo Achiam et al. (2023) - 50.74 34.23
GPT-4o Achiam et al. (2023) - 49.08 32.29
Claude-3 Anthropic (2024a) - 46.14 30.47
OpenMolIns-large (Llama-3.1-8B) 8 43.1 27.22
OpenMolIns-xlarge (Galactica-125M) 0.125 44.48 25.73
Llama3-70B-Instruct (Int4) Dubey et al. (2024) 70 38.54 23.93
OpenMolIns-large (Galactica-125M) 0.125 39.28 23.42
OpenMolIns-medium (Galactica-125M) 0.125 34.54 19.89
GPT-3.5-turbo Achiam et al. (2023) - 28.93 18.58
OpenMolIns-small (Galactica-125M) 0.125 24.17 15.18
Llama3.1-8B-Instruct Dubey et al. (2024) 8 26.26 14.09
Llama3-8B-Instruct Dubey et al. (2024) 8 26.40 13.75
chatglm-9B GLM et al. (2024) 9 18.50 13.13(7)
OpenMolIns-light (Galactica-125M) 0.125 20.95 13.13(6)
OpenMolIns-large (Llama3.2-1B) 1 14.11 8.10
yi-1.5-9B Young et al. (2024) 9 14.10 7.32
Mistral-7B-Instruct-v0.2 Jiang et al. (2023) 7 11.17 4.81
BioT5-base Pei et al. (2023) 0.25 24.19 4.21
MolT5-large Edwards et al. (2022) 0.78 23.11 2.89
Llama-3.1-1B-Instruct Dubey et al. (2024) 1 3.95 1.99
MolT5-base Edwards et al. (2022) 0.25 11.11 1.30(0)
MolT5-small Edwards et al. (2022) 0.08 11.55 1.29(9)
Qwen2-7B-Instruct Yang et al. (2024) 7 0.18 0.15
@@
Then, we present the detailed results of all the subtasks in Table 11, 12, and 13.
@@
Accuracy Similarity Validity Accuracy Similarity Validity Accuracy Similarity Validity
GPT-4o Achiam et al. (2023) 0.6188 0.6782 0.7412 0.7012 0.6038 0.8474 0.7992 0.7225 0.9368
GPT-4-turbo Achiam et al. (2023) 0.699 0.6936 0.7934 0.7244 0.5735 0.906 0.7778 0.7323 0.916
GPT-3.5-turbo Achiam et al. (2023) 0.5832 0.6545 0.798 0.3082 0.7797 0.8468 0.2918 0.6333 0.6822
Claude-3.5 Anthropic (2024b) 0.6832 0.7017 0.4414 0.5414 0.6678 0.796 0.8104 0.731 0.9588
Claude-3 Anthropic (2024a) 0.6766 0.684 0.818 0.5556 0.6408 0.8984 0.655 0.7159 0.9184
Gemini-1.5-pro Deepmind (2024) 0.7058 0.6792 0.8254 0.759 0.5949 0.9158 0.7148 0.7139 0.8684
Llama3-70B-Instruct (Int4) Dubey et al. (2024) 0.5198 0.6801 0.5922 0.6122 0.5637 0.7182 0.5094 0.717 0.6822
Llama3-8B-Instruct Dubey et al. (2024) 0.3914 0.6649 0.5374 0.4348 0.5058 0.57 0.2602 0.6841 0.4838
Llama3.1-8B-Instruct Dubey et al. (2024) 0.2992 0.6088 0.4954 0.4336 0.5257 0.591 0.3401 0.6424 0.5076
Mistral-7B-Instruct-v0.2 Jiang et al. (2023) 0.1868 0.6251 0.376 0.2018 0.3774 0.359 0.0602 0.6227 0.355
Qwen2-7B-Instruct Yang et al. (2024) 0.001 0.2527 0.0036 0.0006 0.4024 0.0012 0.0004 0.2895 0.0068
Yi-1.5-9B Young et al. (2024) 0.1742 0.417 0.4216 0.2858 0.5936 0.4909 0.137 0.4619 0.4368
Chatglm-9B GLM et al. (2024) 0.2932 0.7622 0.5686 0.2956 0.7494 0.6914 0.1498 0.715 0.5084
Llama-3.2-1B-Instruct Dubey et al. (2024) 0.0374 0.5343 0.1982 0.0768 0.575 0.3028 0.0102 0.3671 0.1468
MolT5-small Edwards et al. (2022) 0.122 0.1027 0.449 0.1598 0.1125 0.4504 0.0708 0.1029 0.4876
MolT5-base Edwards et al. (2022) 0.1354 0.1066 0.4686 0.1562 0.1144 0.4472 0.0584 0.1028 0.4426
MolT5-large Edwards et al. (2022) 0.2834 0.1084 0.9282 0.2228 0.1201 0.9198 0.1692 0.0932 0.941
BioT5-base Pei et al. (2023) 0.3462 0.1567 1 0.1668 0.1597 1 0.0684 0.1576 0.9998
OpenMolIns-large (Llama-3.2-1B) 0.1756 0.5676 0.3216 0.1816 0.4963 0.2466 0.0844 0.5415 0.2958
OpenMolIns-large (Llama-3.1-8B) 0.5822 0.6541 0.673 0.5104 0.5074 0.6896 0.544 0.6258 0.84
OpenMolIns-light (Galactica-125M) 0.3786 0.5958 0.3786 0.2062 0.6521 0.7048 0.3102 0.5879 0.6674
OpenMolIns-small (Galactica-125M) 0.3472 0.6172 0.5356 0.3258 0.6025 0.5758 0.2692 0.6181 0.5692
OpenMolIns-medium (Galactica-125M) 0.4736 0.5682 0.7442 0.4886 0.5184 0.7488 0.3282 0.5975 0.6958
OpenMolIns-large (Galactica-125M) 0.5866 0.5876 0.8228 0.6078 0.5577 0.7934 0.3438 0.6491 0.8438
OpenMolIns-xlarge (Galactica-125M) 0.5842 0.5859 0.8438 0.6526 0.5084 0.8286 0.1872 0.6024 0.8538
Table 11: Results on MolEdit. For each task, we highlight the best accuracy and underline the second best accuracy.

### v2

@@ 5 Conclusion
In this study, we introduce TOMG-Bench, the first benchmark designed to assess the capabilities of LLMs in the realm of open-domain molecule generation. Benchmarking 25 LLMs, TOMG-Bench highlights the limitations of existing targeted molecule generation tasks and demonstrates the potential of general LLMs in this domain. Additionally, we propose an instruction tuning dataset, OpenMolIns, to fully explore the potential of LLMs on the TOMG-Bench, enabling Llama-3.1-8B-Instruct to outperform GPT-3.5-turbo. Our contributions not only lie in the development of a novel benchmark for LLMs in molecular discovery but also provide a diverse indication of the LLM evaluation.
Appendix A Comparison with Existing Text-based Molecule Benchmarks or Datasets
Benchmarks or Datasets Task Type # Training Data # Test Data Public Access
PCdes(Zeng et al., 2022) Targeted Generation 10,500 3000 $\times$
ChEBI-20 (Edwards et al., 2021) Targeted Generation 26,407 3,000
PubChem (Liu et al., 2023) Targeted Generation 12,000 2,000
Mol-Instructions (description guided molecule design) (Fang et al., 2024) Targeted Generation 297,319 1000
TOMG-Bench & OpenMolIns Open Generation 1,200,000 45,000
Table 3: Comparison with existing text-based molecule generation benchmarks or datasets.
Table 3 shows the comparison between TOMG-Bench & OpenMolIns and the existing text-based molecule generation benchmarks or datasets. TOMG-Bench introduces the brand-new open-ended molecule generation task for LLMs, which boosts the diversity of text-based molecule generation task. Compared to the existing benchmarks or datasets, OpenMolIns has the largest training set, which could better aligns the molecular space with the natural language space. Meanwhile, TOMG-Bench owns three major tasks and nine subtasks with 45,000 test samples in total, whose test data size is also the largest one among all the existing benchmarks and datasets, enabling more accurate evaluation and comprehensive assessment of LLMs’ performance in text-based molecule generation.
Appendix B Implementation Details
We implement various scripts to facilitate the testing of the aforementioned models. For proprietary models, we adopt the OpenAI API [^1] https://platform.openai.com/docs/ framework. For open-source general LLMs, we utilize both the VLLM [^2] https://docs.vllm.ai framework and the OpenAI framework. For the remaining LLMs, we adopt the Hugging Face transformers library [^3] https://huggingface.co/docs/ for inference.
Furthermore, BioT5 is designed to use SELFIES, an alternative string representation of molecules by using special tokens to ensure valid generations (Krenn et al., 2019), as input instead of SMILES. Consequently, we convert the molecule SMILES strings into SELFIES format on BioT5.
Appendix C Hyper Parameters
Item Value
Generation
temperature 0.75
top_p 0.85
num_beams 1
max_new_tokens 512
Instruction Tuning
batchsize 32
lr 3e-4
cutoff_len 1024
Lora Settings
r 64
$\alpha$ 128
dropout 0.1
Table 4: Hyper-parameters.
Appendix D Data Construction
In this section, we introduce the construction details of TOMG-Bench and OpenMolIns dataset, as well as the prompt templates. Figure 4 demonstrates the overall data construction process for TOMG-Bench as well as the OpenMolIns. Generally, we utilize Zinc250K to construct the test samples in TOMG-Bench, while adopting PubChem for the training samples in OpenMolIns. Notably, when constructing the OpenMolIns, we exclude all the molecules in Zinc250K to avoid data leakage and ensure the novelty score of the generated molecules.
Figure 4: Data construction workflow of TOMG-Bench & OpenMolIns.
D.1 MolEdit
@@
For the MolEdit task, we consider the common operations on modifying functional groups in a given molecule (i.e., add, drop, and substitute), which are simple tasks for human experts but challenging to LLMs. In this case, we further develop three corresponding subtasks: AddComponent, DelComponent, and SubComponent. Prompt templates for MolEdit are shown in Table 5. However, there are different kinds of functional groups, and some functional groups can play an important role in the molecule structure, such as connecting two separate parts of the molecule, which makes them unsuitable for these operations above as these operations will entirely change the structure of the molecule. In this case, we aim to make a slight change in the molecule structure and limit most of the functional groups we choose within the end groups.
@@
Table 5: Prompt Templates for MolEdit.
Table 6 presents the functional groups that are taken into account for AddComponent and DelComponent, along with their respective selection weights. Given that certain functional groups are encountered less frequently in real-world scenarios, we have implemented a weighted random selection process for AddComponent, which ensures that less common functional groups are assigned a lower probability to be chosen.
@@
For SubComponent, our focus is exclusively on end groups for simplification, which include hydroxyl, aldehyde, carboxyl, nitro, halo, nitrile, and thiol, which ensures that the editing operations are confined to substituting the existing functional group with another from this list, thereby maintaining the integrity of the molecule’s overall structure without altering it fundamentally.
D.2 MolOpt
MolOpt is designed to optimize molecular properties through the refinement of molecule structures, is not a brand-new task. Previously, GNN-based methods have been widely adopted in this task, while these methods can only help with one specific subtask at a time. In contrast, TOMG-Bench requires one single LLM to optimize molecules with different metrics and optimization directions. In this work, we specifically focus on enhancing specific characteristics that are crucial for drug discovery and chemical synthesis, including LogP, MR, and QED. The prompt templates for MolOpt are listed in Table 7.
@@ D.2 MolOpt
Table 7: Prompt Templates for MolOpt.
Below are the detailed explanations of the chemical properties adopted in the three subtasks of MolOpt:
@@
D.3 MolCustom
To enable customized design of molecules, we consider three fundamental features for describing the molecule, including atoms, bonds, and functional groups. Given the specified category and number of atoms, bonds, and functional groups, LLMs should generate the molecule as we request. The prompt templates for MolCustom are shown in Table 8. Below, we present the construction details of the three subtasks for MolCustom:
@@ D.3 MolCustom
Table 8: Prompt Templates for MolCustom.
@@
Table 9: Atoms that are considered in AtomNum, as well as their weights to be selected. Higher weights indicate that they are more likely to be selected. For simplification, we only consider normal atoms in the molecule structures.
@@
Table 10: Chemical bonds that are considered in BondNum, as well as their weights to be selected. Higher weights indicate that they are more likely to be selected. For simplification, we only consider normal bonds in the molecule structures.
@@
Table 11: Functional Groups that are considered in FunctionalGroup, as well as their weights to be selected. Higher weights indicate that they are more likely to be selected. For simplification, we only consider normal and edge functional groups in the molecule structures.
@@
Notably, in MolCustom, if not required, LLMs can generate any number of these unspecified atoms, bonds, and functional groups. However, for these specified items, LLMs should strictly follow the requirements.
Appendix E Testing Process
In this section, we present the testing process used to calculate the success rate of each subtask.
Algorithm 1 MolEdit Testing Process
0:  Generated Molecule $m^{g}$, Original Molecule $m^{o}$, Subtask $t$, Group-to-add $\alpha$, Group-to-delete $\Delta$
0:  Pass or Not (Bool) if $t$ is AddComponent then    if $|\alpha|$($m^{g}$) = $|\alpha|$($m^{o})$ + 1 then        return true     else        return false else if $t$ is DelComponent then    if $|\Delta|$($m^{g}$) = $|\Delta|$($m^{o})$ 1 then        return true     else        return false else if $t$ is SubComponent then    if $|\Delta|$($m^{g}$) = $|\Delta|$($m^{o})$ 1 and $|\alpha|$($m^{g}$) = $|\alpha|$($m^{o})$ + 1 then        return true     else        return false
The testing process for MolEdit task is demonstrated in Algorithm 1. However, in this process, we can not guarantee that the modified molecule differs the original molecule only in the specified functional group. The reasons are to follow:
  • The complexity of molecular structures makes it difficult to use regular matching to monitor the exact number of functional groups present. This is because molecules can have a wide variety of bonding patterns and spatial arrangements that are not always amenable to simple, rule-based matching techniques.
  • Modifying functional groups in specific positions can indeed lead to unintended changes in the molecule’s overall structure. This is due to the interconnected nature of atoms within a molecule, where altering one part can have a ripple effect on the stability and geometry of the entire structure.
  • The tools and algorithms currently available may not have the necessary sophistication to predict and control all the possible outcomes of a functional group modification, particularly when dealing with complex molecules or reactions that are not fully understood.
Therefore, the similarity score needs to be considered to ensure the accuracy and quality of the generation.
Algorithm 2 MolOpt Testing Process
0:  : Generated Molecule $m^{g}$, Original Molecule $m^{o}$, Subtask $t$, Optimization Direction Requirement $R$
0:  : Pass or Not (Bool) if $t$ is LogP then    if $LogP$($m^{g}$) $>$ $LogP$($m^{o})$ and R is $\uparrow$ then        return true     else if $LogP$($m^{g}$) $<$ $LogP$($m^{o})$ and R is $\downarrow$ then        return true     else        return false else if $t$ is MR then    if $M\!R$($m^{g}$) $>$ $M\!R$($m^{o})$ and R is $\uparrow$ then        return true     else if $M\!R$($m^{g}$) $<$ $M\!R$($m^{o})$ and R is $\downarrow$ then        return true     else        return false else if $t$ is QED then    if $Q\!E\!D$($m^{g}$) $>$ $Q\!E\!D$($m^{o})$ and R is $\uparrow$ then        return true     else if $Q\!E\!D$($m^{g}$) $<$ $Q\!E\!D$($m^{o})$ and R is $\downarrow$ then        return true     else        return false
For MolOpt, we adopt Algorithm 2 for the testing process. Similarly, merely passing the test does not necessarily mean that the generated molecule follows the principle of minimal change.
Algorithm 3 MolCustom Testing Process
0:  : Generated Molecule $m^{g}$, Subtask $t$, Atom List $A$, Bond List $B$, Functional Group List $G$, Requirements $R$
0:  : Pass or Not (Bool) flag = trueif $t$ is AtomNum then    for $atom$ in $A$ do         if $|atom|$($m^{g}$) $\neq$ $R$[$atom$then            flag = false else if $t$ is BondNum then    for $bond$ in $B$ do         if $|bond|$($m^{g}$) $\neq$ $R$[$bond$then            flag = false else if $t$ is FunctionalGroup then    for $group$ in $G$ do         if $|group|$($m^{g}$) $\neq$ $R$[$group$then            flag = false return flag
For MolCustom, the testing process is shown in Algorithm 3, which test whether the specified number of the atom, bond, or functional group satisfies the requirements.
Appendix F Detailed Results
@@
In this section, we first show the leaderboard of TOMG-Bench in Table 12, where Claude-3.5 achieves first place with an average weighted success rate of 35.92%. Notably, via instruction tuning on our OpenMolIns dataset, Llama-3.1-8B achieves 6th place, which outperforms all the existing open-source general LLMs.
Model #Parameters (B) $\overline{S\!R}$ (%) $\overline{W\!S\!R}(\%)$
Claude-3.5 (Anthropic, 2024b) - 51.10 35.92
Gemini-1.5-pro (Deepmind, 2024) - 52.25 34.80
GPT-4-turbo (OpenAI, 2023) - 50.74 34.23
GPT-4o (OpenAI, 2023) - 49.08 32.29
Claude-3 (Anthropic, 2024a) - 46.14 30.47
Llama-3.1-8B (OpenMolIns-large) 8 43.1 27.22
Galactica-125M (OpenMolIns-xlarge) 0.125 44.48 25.73
Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 70 38.54 23.93
Galactica-125M (OpenMolIns-large) 0.125 39.28 23.42
Galactica-125M (OpenMolIns-medium) 0.125 34.54 19.89
GPT-3.5-turbo (OpenAI, 2023) - 28.93 18.58
Galactica-125M (OpenMolIns-small) 0.125 24.17 15.18
Llama3.1-8B-Instruct (Dubey et al., 2024) 8 26.26 14.09
Llama3-8B-Instruct (Dubey et al., 2024) 8 26.40 13.75
chatglm-9B (GLM et al., 2024) 9 18.50 13.13(7)
Galactica-125M (OpenMolIns-light) 0.125 20.95 13.13(6)
Llama3.2-1B (OpenMolIns-large) 1 14.11 8.10
yi-1.5-9B (Young et al., 2024) 9 14.10 7.32
Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 7 11.17 4.81
BioT5-base (Pei et al., 2023) 0.25 24.19 4.21
MolT5-large (Edwards et al., 2022) 0.78 23.11 2.89
Llama-3.1-1B-Instruct (Dubey et al., 2024) 1 3.95 1.99
MolT5-base (Edwards et al., 2022) 0.25 11.11 1.30(0)
MolT5-small (Edwards et al., 2022) 0.08 11.55 1.29(9)
Qwen2-7B-Instruct (Yang et al., 2024) 7 0.18 0.15
@@
Then, we compare several representative LLMs from different model categories via radar maps, which are shown in Figure 5. Obviously, it can be concluded that Claude-3.5 outperforms almost all the other LLMs in these subtasks, except for the DelComponent task and the BondNum task. In DelComponent task, GPT-4o demonstrates better performance than Claude-3.5, while Llama-3.1-8B (OpenMolIns-large) achieves the best performance in the BondNum task.
Meanwhile, these radar maps also imply the difficulties of the subtasks. For example, for majority of the LLMs, they are struggling with the BondNum task. Even the most advanced proprietary LLMs like Claude-3.5 failed to achieve an ideal performance.
Moreover, in the MolOpt task, QED, which assesses the potential of a molecule to become a drug for curing diseases, is generally considered the most challenging property for molecule optimization. In contrast, LogP and MR can be more directly inferred from the molecule’s structure, thus making them more straightforward to optimize.
Figure 5: The radar maps compare the performance ($\overline{W\!S\!R}$) of some representative LLMs from different model categories in each subtask of TOMG-Bench.
Finally, we present the detailed experimental results of all the subtasks in Table 13, 14, and 15.
@@
SR Similarity WSR Validity SR Similarity WSR Validity SR Similarity WSR Validity
GPT-4o (OpenAI, 2023) 0.6188 0.6782 0.4197 0.7412 0.7012 0.6038 0.4234 0.8474 0.7992 0.7225 0.5774 0.9368
GPT-4-turbo (OpenAI, 2023) 0.6990 0.6936 0.4848 0.7934 0.7244 0.5735 0.4154 0.9060 0.7778 0.7323 0.5696 0.9160
GPT-3.5-turbo (OpenAI, 2023) 0.5832 0.6545 0.3817 0.7980 0.3082 0.7797 0.2403 0.8468 0.2918 0.6333 0.1848 0.6822
Claude-3.5 (Anthropic, 2024b) 0.6832 0.7017 0.4794 0.4414 0.5414 0.6678 0.3615 0.7960 0.8104 0.7310 0.5924 0.9588
Claude-3 (Anthropic, 2024a) 0.6766 0.6840 0.4628 0.8180 0.5556 0.6408 0.3560 0.8984 0.6550 0.7159 0.4689 0.9184
Gemini-1.5-pro (Deepmind, 2024) 0.7058 0.6792 0.4794 0.8254 0.7590 0.5949 0.4515 0.9158 0.7148 0.7139 0.5103 0.8684
Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 0.5198 0.6801 0.3535 0.5922 0.6122 0.5637 0.3451 0.7182 0.5094 0.7170 0.3652 0.6822
Llama3-8B-Instruct (Dubey et al., 2024) 0.3914 0.6649 0.2602 0.5374 0.4348 0.5058 0.2199 0.5700 0.2602 0.6841 0.1780 0.4838
Llama3.1-8B-Instruct (Dubey et al., 2024) 0.2992 0.6088 0.1822 0.4954 0.4336 0.5257 0.2279 0.5910 0.3401 0.6424 0.2185 0.5076
Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 0.1868 0.6251 0.1168 0.3760 0.2018 0.3774 0.0762 0.3590 0.0602 0.6227 0.0375 0.3550
Qwen2-7B-Instruct (Yang et al., 2024) 0.0010 0.2527 0.0003 0.0036 0.0006 0.4024 0.0002 0.0012 0.0004 0.2895 0.0001 0.0068
Yi-1.5-9B (Young et al., 2024) 0.1742 0.4170 0.0726 0.4216 0.2858 0.5936 0.1697 0.4909 0.1370 0.4619 0.0633 0.4368
Chatglm-9B (GLM et al., 2024) 0.2932 0.7622 0.2235 0.5686 0.2956 0.7494 0.2215 0.6914 0.1498 0.7150 0.1071 0.5084
Llama-3.2-1B-Instruct (Dubey et al., 2024) 0.0374 0.5343 0.0200 0.1982 0.0768 0.5750 0.0442 0.3028 0.0102 0.3671 0.0037 0.1468
MolT5-small (Edwards et al., 2022) 0.1220 0.1027 0.0125 0.4490 0.1598 0.1125 0.0180 0.4504 0.0708 0.1029 0.0073 0.4876
MolT5-base (Edwards et al., 2022) 0.1354 0.1066 0.0144 0.4686 0.1562 0.1144 0.0179 0.4472 0.0584 0.1028 0.0060 0.4426
MolT5-large (Edwards et al., 2022) 0.2834 0.1084 0.0307 0.9282 0.2228 0.1201 0.0268 0.9198 0.1692 0.0932 0.0158 0.9410
BioT5-base (Pei et al., 2023) 0.3462 0.1567 0.0542 1.0000 0.1668 0.1597 0.0266 1.0000 0.0684 0.1576 0.0108 0.9998
Llama-3.2-1B (OpenMolIns-large) 0.1756 0.5676 0.0997 0.3216 0.1816 0.4963 0.0901 0.2466 0.0844 0.5415 0.0457 0.2958
Llama-3.1-8B (OpenMolIns-large) 0.5822 0.6541 0.3808 0.6730 0.5104 0.5074 0.2590 0.6896 0.5440 0.6258 0.3404 0.8400
Galactica-125M (OpenMolIns-light) 0.3786 0.5958 0.2256 0.6842 0.2062 0.6521 0.1345 0.7048 0.3102 0.5879 0.1824 0.6674
Galactica-125M (OpenMolIns-small) 0.3472 0.6172 0.2143 0.5356 0.3258 0.6025 0.1963 0.5758 0.2692 0.6181 0.1664 0.5692
Galactica-125M (OpenMolIns-medium) 0.4736 0.5682 0.2691 0.7442 0.4886 0.5184 0.2533 0.7488 0.3282 0.5975 0.1961 0.6958
Galactica-125M (OpenMolIns-large) 0.5866 0.5876 0.3447 0.8228 0.6078 0.5577 0.3390 0.7934 0.3438 0.6491 0.2232 0.8438
Galactica-125M (OpenMolIns-xlarge) 0.5842 0.5859 0.3423 0.8438 0.6526 0.5084 0.3318 0.8286 0.1872 0.6024 0.1128 0.8538
Table 13: Detailed results on MolEdit. For each task, we highlight and underline the best success rate (SR), as well as the weighted success rate (WSR).